OpenStudy (mathlegend):

Word Problem: You have started a business making papier mache sculptures. The cost to make a sculpture is $.75. Your sculptures sell for $14.50 each at a craft store. You receive 50% of the selling price. Each sculpture takes about 2 hours to complete. If you spend 14 hours per week making sculptures, about how many weeks will you work to earn a profit of $360?

OpenStudy (mathlegend):

My answer as a decimal was 7.1

OpenStudy (mathlegend):

Is this right?

OpenStudy (mathlegend):

@hartnn are you good at word problems?

hartnn (hartnn):

i can try, wait.

hartnn (hartnn):

i think its not given how many hours it requires to sell a sculpture? do i assume in 14 hours 1 sculpture sold ?

OpenStudy (mathlegend):

2 hours to complete 1 sculpture

OpenStudy (mathlegend):

I suppose every single one makes a profit

hartnn (hartnn):

so 7 profits per day, so 7* 6.5 =45.5 $ profit per day, isn't it ?

hartnn (hartnn):

360/45.5 which is 7.912 days or 1 week 1 day

OpenStudy (mathlegend):

Per week I got $50.75

OpenStudy (mathlegend):

I got 50% of 14.50 = $7.25

hartnn (hartnn):

per day ?

OpenStudy (mathlegend):

Per each individual item

hartnn (hartnn):

ok, thats what he gets , but profit, will be 7.25 - 0.75 = 6.5 isn't it ?? and that too in 2 hours as u stated ....ok?

OpenStudy (mathlegend):

ok

hartnn (hartnn):

so per day, it would be 6.5*7 right ?

OpenStudy (mathlegend):

6.5 Per sculpture right?

hartnn (hartnn):

yes.

OpenStudy (mathlegend):

ok

OpenStudy (mathlegend):

Okay so 45.5 * x = 360?

hartnn (hartnn):

thats what i got...

OpenStudy (mathlegend):

ok

OpenStudy (mathlegend):

so do I round 7.9 to 8?

hartnn (hartnn):

yup, unless 'u receive 50% of selling price' means thats the profit.... as i have calculated frofit as 7.25-0.75.....

OpenStudy (mathlegend):

Thanks

hartnn (hartnn):

welcome :)